Logiweb bibliography of prop

Up Help

[0] prop

[1] base

[2] check

The pyk compiler, version 0.grue.20050502+ by Klaus Grue GRD-2005-06-02.UTC:08:43:13.701947 = MJD-53523.TAI:08:43:45.701947 = LGT-4624418625701947e-6